What total volume,in litre at 600 degree celcius and 1 atm ,could be formed by the decomposition of 16gm of NH4NO3     
Reaction: 2NH4NO3--> 2N2+O2+4H2O(g)

Grade:12th pass

2 Answers

Arun
4 years ago
Ammonium Nitrate is decomposed as,
NH4NO3 -> N2O + 2H2O,
According to stoichiometric coeffiecients, 1 mole of NH4NO3 dissociates to form 1 mole of nitrous oxide and 2 moles of water, but in the question we have 16g of NH4NO3,
So the number of moles of NH4NO3 = 16/mass of 1 mole of NH4NO3 = 16/80 = 0.2 mole,
Use the ideal gas equation PV = nRT,
V = (0.2)(0.082)(873)/1 = 14.3172 L
 
Khimraj
4 years ago

MOLES OF NH4NO3 = 16/80= 1/5

NOW THE REACTION IS 

1/5 NH4NO3 ------>  1/5 N2 + 1/10 O2 + 2/5 H2O

MOLES N = 1/5 + 1/10 = 3/10

NOW BY GAS EQ. 

PV=NRT 

V=21.5L